IF N= -3 AND X =(N_5), WHAT IS THE VALUE OF X?
Answer by venugopalramana(3286)

IF N= -3 AND X =(N_5), WHAT IS THE VALUE OF X?
it is assumed that X =(N_5), is a typing mistake and it really means N-5..
we have to put N=-3 in the the given equation for X=N-5=-3-5=-8